# A Review of Traffic Wave Suppression Strategies: Variable Speed Limit vs. Jam-Absorption Driving

### v1

@@ Abstract
The main form of freeway traffic congestion is the familiar stop-and-go wave, characterized by wide moving jams that propagate indefinitely upstream provided enough traffic demand. They cause severe, long-lasting adverse effects, such as reduced traffic efficiency, increased driving risks, and higher vehicle emissions. This underscores the crucial importance of artificial intervention in the propagation of stop-and-go waves. Over the past two decades, two prominent strategies for stop-and-go wave suppression have emerged: variable speed limit (VSL) and jam-absorption driving (JAD). Although they share similar research motivations, objectives, and theoretical foundations, the development of these strategies has remained relatively disconnected. To synthesize fragmented advances and drive the field forward, this paper first provides a comprehensive review of the achievements in the stop-and-go wave suppression-oriented VSL and JAD, respectively. It then focuses on bridging the two areas and identifying research opportunities from the following perspectives: fundamental diagrams, traffic dynamics modeling, traffic state estimation and prediction, stochasticity, scenarios for strategy validation, and field tests and practical deployment. We expect that through this review, one area can effectively address its limitations by identifying and leveraging the strengths of the other, thus promoting the overall research goal of freeway stop-and-go wave suppression.
@@
Intelligent Transportation Systems, traffic congestion, traffic waves, travel time, emission, vehicle dynamics
@@ I-A Stop-and-Go Waves on Freeways
As shown in Figure 1, stop-and-go waves (also referred to as wide moving jams, shock waves, or traffic oscillations) are a common form of traffic congestion on freeways worldwide. In the time-space diagrams of traffic speed, a stop-and-go wave appears as a striped low-speed region propagating upstream, against the direction of traffic flow. Physically, on the freeway, a stop-and-go wave is essentially a backward-moving queue with an approximately constant queue length. As vehicles traverse these queues, they typically undergo a sequence of deceleration, low-speed cruising or even standstill, followed by acceleration.
@@ I-C Motivation
  • Among the selected VSL and JAD papers identified as representative works, none of the VSL papers cite JAD studies, whereas JAD papers tend to reference VSL works, albeit not comprehensively (Figure 2).
@@ I-D Contribution
To fulfill the above motivations, this paper systematically reviews the current progress in the suppression of freeway stop-and-go waves exclusively through two techniques: VSL and JAD. We demonstrate that, despite the many similarities (Table I), the two areas have developed distinct research focuses. For instance, VSL studies are strong in theory-based modeling but often lack a thorough consideration of secondary waves. In contrast, JAD studies are focused more on the microscopic perspective but lack a solid theoretical foundation and field test. By leveraging the strengths of one area, the other can effectively address its limitations, thus promoting the overall research goal of stop-and-go wave suppression.
Moreover, we identify key research gaps in both areas and highlight future opportunities, drawing not only on foundational traffic flow theories and advanced artificial intelligence (AI) techniques but also by bridging the development paths of VSL and JAD in a unified framework. In searching and reviewing related papers, we did find that some studies were not developing along sound directions. We, therefore, hope this review not only inspires future work but also helps avoid inappropriate assumptions and methods in the future. Meanwhile, this review would also be beneficial for researchers in neighboring fields, such as connected and automated vehicle (CAV)-based traffic control, ramp metering, and even eco-driving and traffic signal control at intersections.
@@ I-E Scope
In addition, although some JAD studies position themselves in the framework of CAV technologies, we argue that we may not have to rely on the technologies when implementing JAD. In Section III-F, we will show that the swerving behavior of a police car is practically observed on California freeways to suppress traffic at high speeds. This observation greatly expands the potential for practical applications of JAD, especially at present, and encourages us to position JAD beyond CAV-based strategies.
@@ I-F Organization
The rest of the paper is organized as follows. Section II provides a detailed overview of the developments in VSL and JAD strategies, introducing, organizing, and comparing representative studies. Section III bridges the two areas and identifies future research opportunities from the following perspectives: fundamental diagrams (FDs), traffic dynamics modeling, traffic state estimation and prediction, stochasticity, scenarios for strategy validation, and field tests and practical deployment. Section IV concludes the review by summarizing the main findings and key insights.
@@
This section presents the development in freeway stop-and-go wave suppression through VSL and JAD. Key features of VSL and JAD studies, selected based on our review, are summarized in Table III, the details of which will not be reiterated individually in the main text unless they are particularly important and warrant specific discussion.
@@ II-A Variable Speed Limit
Primarily, VSL strategies have been developed along two main threads: model predictive control (MPC)-based strategies and kinematic wave theory (KWT)-based strategies. In the early stage, the MPC was built upon METANET [39], which is a second-order cell-based macroscopic traffic flow model [61, 15, 62]. The strategies divide a freeway stretch into a series of cells (i.e., segments). Each cell, approximately 300 to 1000 meters in length, corresponds to a VSL. The MPC-based strategy optimally coordinates VSL at different segments with the objective of minimizing total travel time. It is known that the effectiveness of MPC is directly determined by the accuracy of the predictive model used, turning out that the representativeness of METANET is the key to the effectiveness of the MPC-based strategies. Initially, [39] tested the METANET-based MPC solely within a METANET-based simulation, which is equivalent to assuming perfect prediction. To relax the assumption, [40] conducted a more rigorous and practical validation using the microscopic traffic flow simulation software Paramics 5.1, a commercial tool developed by Quadstone. Further, to address the high computational burden associated with nonlinear and non-convex MPC, [41] proposed a linear MPC-based strategy with a modified first-order cell transmission model (CTM), achieving similar optimization results with significantly improved efficiency. More recently, [42] introduced a linear Lagrangian MPC, which utilizes average vehicle speeds as decision variables rather than flow, as in previous MPC-based strategies. This modification bridges traffic-oriented control and vehicle-oriented control, enhancing the flexibility of the overall framework of the stop-and-go wave suppression.
@@
Another thread is KWT-based VSL strategies. [44] proposed the seminal algorithm (named SPECIALIST) by conducting a theoretical analysis using KWT on an FD with a capacity drop (i.e., $\lambda$shape FD). The opportunity to suppress a stop-and-go wave arises from the empirical observation of an outflow reduction of up to 30% associated with a stop-and-go wave [63]. Later, in 2010, [21] conducted a field test of SPECIALIST to suppress stop-and-go waves over a 14 km stretch of the Dutch A12 freeway, which, to date, may be the only real-world test of such stop-and-go wave suppression strategies. Although the effectiveness of the algorithm was well demonstrated (Figure 3(a)(b)), it was found that the algorithm was frequently activated for other types of jams rather than stop-and-go waves. Notably, it was reported that secondary waves were triggered by the newly imposed lower speed limit (Figure 3(c)(d)). On this basis, [45] introduced ramp metering to the SPECIALIST framework and proposed a coordinated strategy with mainstream control and ramp metering. To take advantage of CAVs, [46] introduced a bi-level framework to combine VSL and CAVs and enhance the suppression (actually earlier than [42], where a similar bi-level framework was adopted). To deliver VSL directly to vehicles, the latest work reformulated a $\lambda$shape FD into a triangular form and proposed a Lagrangian strategy capable of generating VSL suitable for implementation in connected vehicles [47]. It is important to note that the theory behind SPECIALIST and its variants holds only for isolated stop-and-go waves with high-density free-flow conditions upstream and downstream (such as the one with label “A jam wave” in Figure 1(d)), which are different from most waves presented in Figure 1. In contrast, although MPC-based strategies were also proposed to dissipate the same type of waves, their underlying mechanisms and optimization objectives suggest that they may be effective for a broader range of wave types, provided that the control resolution is increased. As we will show in what follows, JAD does not specifically target this type of waves.
@@
Figure 3: Examples of the field test results in [21]. (a)(b) A successful resolved stop-and-go wave: speed (km/h) and flow (veh/h/lane); (c)(d) A case where secondary waves were triggered: speed (km/h) and flow (veh/h/lane).
@@
Even though secondary waves were observed in the real world as early as 2010 [21], subsequent studies have not paid much attention to this issue. This contrasts sharply with JAD strategies, which have extensively discussed the secondary wave issue, as we will show in Section II-B. Another reason why the secondary wave issue has not received much attention may lie in the selection of simulation models. Some microscopic models and macroscopic models may be too stable to replicate traffic breakdowns caused by, for instance, the imperfect driving behavior of human drivers.
@@ II-B Jam-Absorption Driving
The concept of JAD was initially introduced in a manner similar to a popular science article, as discussed by Beaty on a website [65, 66]. It was not until 2013 that [51] sought to systematically examine and analyze its potential in a scientifically rigorous manner. From the very beginning, the research questions were framed around two key questions: (1) Can JAD suppress stop-and-go waves? (2) Will secondary waves be triggered? Centered around these two questions, [51] conducted a theoretical analysis based on KWT, focusing more on the impact of FD shapes (i.e., concave, convex, and triangular FDs) rather than on a capacity drop, which was the main focus of KWT-based VSL studies [44]. It concluded that JAD is effective only when the inflow is lower than the capacity. Unfortunately, this condition does not support the self-sustainability of a stop-and-go wave, as demonstrated by the empirical observations in Figure 1. We will discuss this in detail in Section III-A. Later, [52, 53] extended the analysis from a microscopic perspective by employing the Helly CF model and the intelligent driver model (IDM). The explicit formulations of vehicle dynamics in these CF models make the analyses traceable. However, it is important to be aware that the conclusions rely on the strong assumptions inherent in these models.
@@
A closer look at the JAD-involved simulation reveals potential opportunities for wave suppression: as shown in Figure 5, compared to the queue discharge rate at A, JAD results in a higher flow at B, even though the speed at B cannot exceed that at A due to the constraint imposed by the vehicles at A. It should also be noted that the resulting high flow may be unstable and could lead to further breakdowns.
@@
Up until now, according to the latest publication [57], the issue of whether secondary waves could be triggered remains unresolved. The newly triggered secondary waves may largely contribute to increased travel time when absorbing dense stop-and-go waves [54]. Drawing on insights from VSL strategies, we recommend incorporating the empirically observed outflow reduction caused by stop-and-go waves and the potential benefits of mitigating this phenomenon into the modeling process, as it may serve as a key remedy for improving traffic efficiency. As shown in [13], the queue discharge rate reduction is positively and linearly related to the average speed of vehicles within a moving queue, highlighting the potential benefit of suppressing a wave, especially when the speed within the queue is low.
@@
It is worth noting that we have observed a growing trend in which a platoon of vehicles has been employed and controlled to implement JAD for suppressing stop-and-go waves [58, 59, 60, 56]. However, this approach seems to deviate from the original motivation behind JAD strategies, which is to suppress stop-and-go waves using a single (or very few) vehicle(s). We should be clearly aware that if we have the capability to stably and precisely control a vehicle platoon, the platoon-control-based JAD strategies should be compared with high CAV penetration-based speed homogenization strategies [29, 30, 31, 32, 33, 34], rather than with the single-vehicle-based JAD strategies. In Section III-F, we will illustrate that single-vehicle-based JAD is actually quite feasible in practice.
@@ III-A Fundamental Diagrams
Many existing strategies and analyses are based on the FDs or their Lagrangian variants (e.g., speed-spacing relationship). It can be asserted that the shape of the assumed FDs directly influences both the formulation of a strategy and its effectiveness. As shown in [68], the shape of an FD is closely related to its spatial relation to a bottleneck. Specifically, an FD observed immediately downstream of a bottleneck typically exhibits a clear capacity drop, whereas upstream of the bottleneck, the FD tends to approximate a triangular shape (Figure 6). Therefore, when assuming an FD, it is crucial to first determine the location or the type of the stop-and-go waves.
@@
For the SPECIALIST, it targets isolated stop-and-go waves (such as the one with the label “A jam wave” in Figure 1(d)). Since the upstream and downstream traffic of such waves are in high-density free-flow conditions, the wave can be regarded as a bottleneck, and the associated FD may exhibit a capacity drop (Figure 7(b)). This assumed $\lambda$shape FD effectively resolves the contradiction between the widely observed constant length of the stop-and-go wave and its shrinking after imposing a lower speed limit. More specifically, to be consistent with real-world observations (i.e., a constant length of the moving queue), the inflow and outflow of the wave must be equal (i.e., at states 1 and 6 in Figures 7(a) and (b)). Without a capacity drop (i.e., no state 5), traffic will return from state 4 to state 1 (illustrated by the blue line and arrow in Figures 7(a) and (b), respectively), leading to the formation of the secondary wave at state 4. More importantly, this wave at state 4 will not shrink and may even continue to grow, although the density at state 4 is lower than the wave density at state 2. It is noted that state 5 is highly unstable due to high flow [69]. Therefore, even though a stop-and-go wave is suppressed, as shown in Figure 7(a), it remains uncertain how long the high-flow region (state 5) can be sustained.
@@
In contrast, JAD strategies target more broadly defined stop-and-go waves, rather than focusing solely on isolated waves. When multiple waves occur, as shown in the time-space diagrams in Figure 1, the traffic across the multiple and periodic stop-and-go waves is all in the congestion regime, and the corresponding FD is a triangle with no distinct capacity drop. Many JAD strategies do not make this distinction clearly, as the original JAD idea is simply to perform “slow-in” and fast-out” actions. For example, state b in Figure 7(c) is still self-sustained [54]. To achieve a self-shrinking JAD-impacted region with the triangular FD, [51] concluded that an effective JAD requires the inflow to be lower than the capacity in the triangular FD (e.g., inflow at state d in Figure 7(d)). It ensures the shrinking of the wave after JAD, but the condition implies that the wave, where JAD can take effect, can also shrink on its own (although it may take time), regardless of whether JAD control is applied (see the green line in Figure 7(c)). More rigorous efforts are called for to establish a robust theoretical foundation grounded in empirical evidence.
Another two assumptions about FDs are also worth mentioning. The first is the use of three-phase FDs in [48, 50], which leads to different conclusions compared to the ones based on the conventional two-phase FDs [44]. But we also remind that the TPTT may be explainable by the triangular FD-based KWT and lane-changing (LC) behavior [68]. The second one incorporates the state that is higher than the triangular-shaped states (i.e., state 5 in Figure 7(b)) into the triangle [60, 47] or modifies the shape of the triangular FD to accommodate the higher capacity [41, 70], thereby achieving the goal of simplifying the modeling.
@@
III-B Traffic Dynamics Modeling

### v2

# A Review of Stop-and-Go Traffic Wave Suppression Strategies: Variable Speed Limit vs. Jam-Absorption Driving

@@ Abstract
The main form of freeway traffic congestion is the familiar stop-and-go wave, characterized by wide moving jams that propagate indefinitely upstream provided enough traffic demand. They cause severe, long-lasting adverse effects, such as reduced traffic efficiency, increased driving risks, and higher vehicle emissions. This underscores the crucial importance of artificial intervention in the propagation of stop-and-go waves. Over the past two decades, two prominent strategies for stop-and-go wave suppression have emerged: variable speed limit (VSL) and jam-absorption driving (JAD). Although they share similar research motivations, objectives, and theoretical foundations, the development of these strategies has remained relatively disconnected. To synthesize fragmented advances and drive the field forward, this paper first provides a comprehensive review of the achievements in the stop-and-go wave suppression-oriented VSL and JAD, respectively. It then focuses on bridging the two areas and identifying research opportunities from the following perspectives: fundamental diagrams, secondary waves, generalizability, traffic state estimation and prediction, robustness to randomness, scenarios for strategy validation, and field tests and practical deployment. We expect that through this review, one area can effectively address its limitations by identifying and leveraging the strengths of the other, thus promoting the overall research goal of freeway stop-and-go wave suppression.
@@
Traffic congestion, traffic control, travel time, vehicle dynamics, mobility, artificial intelligence
@@ I-A Stop-and-Go Waves on Freeways
As shown in Figure 1, stop-and-go waves (also referred to as wide moving jams or traffic oscillations) are a common form of traffic congestion on freeways worldwide. In the time-space diagrams of traffic speed, a stop-and-go wave appears as a striped low-speed region propagating upstream, against the direction of traffic flow. Physically, on the freeway, a stop-and-go wave is essentially a backward-moving queue with an approximately constant queue length. As vehicles traverse these queues, they typically undergo a sequence of deceleration, low-speed cruising or even standstill, followed by acceleration.
@@ I-B Stop-and-Go Wave Suppression Strategies
  • 1
    $\dagger$ VSL acts on the traffic flow from outside the flow itself.
  • 2
    $\ddagger$ The JAD vehicle acts within the traffic flow as an integral part of it.
@@ I-C Motivation
  • Among the selected VSL and JAD papers identified as representative works, none of the VSL papers cite JAD studies, even though, as we will show later, a trend toward incorporating connected and automated vehicles (CAVs) into VSL is observed. In contrast, JAD papers tend to reference VSL works, albeit not comprehensively (Figure 2).
@@ I-D Contribution
To fulfill the above motivations, this paper systematically reviews the current progress in strategies proposed to suppress freeway stop-and-go waves, exclusively through two techniques: VSL and JAD. We demonstrate that, despite the many similarities (Table I), the two areas have developed distinct research focuses. For instance, VSL studies are strong in theory-based modeling but often lack a thorough consideration of secondary waves that may be triggered by the reduction of speed. In contrast, JAD studies are focused more on the microscopic perspective but lack a solid theoretical foundation and field test. By leveraging the strengths of one area, the other can effectively address its limitations, thus promoting the overall research goal of stop-and-go wave suppression.
Moreover, we identify key research gaps in both areas and highlight future opportunities, drawing not only on foundational traffic flow theories and advanced artificial intelligence (AI) techniques but also by bridging the development paths of VSL and JAD in a unified framework. In searching and reviewing related papers, we did find that some studies were not developing along sound directions. We, therefore, hope this review not only inspires future work but also helps avoid inappropriate assumptions and methods in the future. Meanwhile, this review would also be beneficial for researchers in neighboring fields, such as CAV-based traffic control, ramp metering, and even eco-driving and traffic signal control at intersections.
@@ I-E Scope
In addition, although some JAD studies position themselves in the framework of CAV technologies, we argue that we may not have to rely on the technologies when implementing JAD. In Section III-G, we will show that the swerving behavior of a police car is practically observed on California freeways to suppress traffic at high speeds. This observation greatly expands the potential for practical applications of JAD, especially at present, and encourages us to position JAD beyond CAV-based strategies.
@@ I-F Organization
The rest of the paper is organized as follows. Section II provides a detailed overview of the developments in VSL and JAD strategies, introducing, organizing, and comparing representative studies. Section III bridges the two areas and identifies future research opportunities from the following perspectives: fundamental diagrams (FDs), secondary waves, generalizability of strategies, traffic state estimation and prediction, stochasticity, scenarios for strategy validation, and field tests and practical deployment. Section IV concludes the review by summarizing the main findings and key insights.
@@
This section presents the development in freeway stop-and-go wave suppression strategies through VSL and JAD. Key features of VSL and JAD studies, selected based on our review, are summarized in Table III, the details of which will not be reiterated individually in the main text unless they are particularly important and warrant specific discussion.
@@ II-A Variable Speed Limit
Primarily, VSL strategies have been developed along two main threads: model predictive control (MPC)-based strategies and kinematic wave theory (KWT)-based strategies. In the early stage, the MPC was built upon METANET [39], which is a second-order, deterministic, cell-based macroscopic traffic flow model [61, 15, 62]. The strategies divide a freeway stretch into a series of cells (i.e., segments). Each cell, approximately 300 to 1000 meters in length, corresponds to a VSL. The MPC-based strategy optimally coordinates VSL at different segments with the objective of minimizing total travel time. It is known that the effectiveness of MPC is directly determined by the accuracy of the predictive model used, turning out that the representativeness of METANET is the key to the effectiveness of the MPC-based strategies. Initially, [39] tested the METANET-based MPC solely within a METANET-based simulation, which is equivalent to assuming perfect prediction. To relax the assumption, [40] conducted a more rigorous and practical validation using the microscopic traffic flow simulation software Paramics 5.1, a commercial tool developed by Quadstone. Further, to address the high computational burden associated with nonlinear and non-convex MPC, [41] proposed a linear MPC-based strategy with a modified first-order cell transmission model (CTM), achieving similar optimization results with significantly improved efficiency. More recently, [42] introduced a linear Lagrangian MPC, which utilizes average vehicle speeds as decision variables rather than flow, as in previous MPC-based strategies. This modification bridges traffic-oriented control and vehicle-oriented control, enhancing the flexibility of the overall framework of the stop-and-go wave suppression.
@@
Another thread is KWT-based VSL strategies. [44] proposed the seminal algorithm (named SPECIALIST) by conducting a theoretical analysis using KWT on an FD with a capacity drop (i.e., $\lambda$shape FD). The opportunity to suppress a stop-and-go wave arises from the empirical observation of an outflow reduction of up to 30% associated with a stop-and-go wave [63]. Later, in 2010, [21] conducted a field test of SPECIALIST to suppress stop-and-go waves over a 14 km stretch of the Dutch A12 freeway, which, to date, may be the only real-world test of such stop-and-go wave suppression strategies. Although the effectiveness of the algorithm was well demonstrated (Figure 3(a)(b)), it was found that the algorithm was frequently activated for other types of jams rather than stop-and-go waves. Notably, it was reported that secondary waves were triggered by the newly imposed lower speed limit (Figure 3(c)(d)). [45] introduced ramp metering to the SPECIALIST framework and proposed a coordinated strategy with mainstream control and ramp metering. To take advantage of CAVs, [46] introduced a bi-level framework to combine VSL and CAVs and enhance the suppression (actually earlier than [42], where a similar bi-level framework was adopted). To deliver VSL directly to vehicles, the latest work reformulated a $\lambda$shape FD into a triangular form and proposed a Lagrangian strategy capable of generating VSL suitable for implementation in connected vehicles [47]. It is important to note that SPECIALIST and its variants are built upon FDs that exhibit capacity drop, meaning they are only applicable to stop-and-go waves where capacity drop occurs. As we will elaborate in Section III-A, only isolated stop-and-go waves—with high-density free-flow conditions both upstream and downstream—may trigger capacity drop. An example is the wave labeled “A jam wave” in Figure 1(d), which differs from most of the waves shown in Figure 1. In contrast, although MPC-based strategies were also proposed to dissipate the same type of waves, their underlying mechanisms and optimization objectives (i.e., minimizing total time spent) suggest that they may be effective for a broader range of wave types, provided that the control resolution is increased. As we will show in what follows, JAD does not specifically target this type of waves.
@@
Figure 3: Examples of the field test results in [21]. (a)(b) A successfully resolved stop-and-go wave: speed (km/h) and flow (veh/h/lane); (c)(d) A case where secondary waves were triggered: speed (km/h) and flow (veh/h/lane).
@@
Even though secondary waves were observed in the real world as early as 2010 [21], subsequent VSL studies have not paid much attention to this issue. This contrasts sharply with JAD strategies, which have extensively discussed the secondary wave issue, as we will show in Section II-B. Another reason why the secondary wave issue has not received much attention may lie in the selection of simulation models. Some microscopic models and macroscopic models may be too stable to replicate traffic breakdowns caused by, for instance, the imperfect driving behavior of human drivers.
@@ II-B Jam-Absorption Driving
The concept of JAD was initially introduced in a manner similar to a popular science article, as discussed by Beaty on a website [65, 66]. It was not until 2013 that [51] sought to systematically examine and analyze its potential in a scientifically rigorous manner. From the very beginning, the research questions were framed around two key questions: (1) Can JAD suppress stop-and-go waves? (2) Will secondary waves be triggered? Centered around these two questions, [51] conducted a theoretical analysis based on KWT, focusing more on the impact of FD shapes (i.e., concave, convex, and triangular FDs) rather than on a capacity drop, which was the main focus of KWT-based VSL studies [44]. It concluded that JAD is effective only when the inflow is lower than the capacity. Unfortunately, this condition does not support the self-sustainability of a stop-and-go wave that has been well demonstrated by the empirical observations in Figure 1. We will discuss this in detail in Section III-A. Later, [52, 53] extended the analysis from a microscopic perspective by employing the Helly CF model and the intelligent driver model (IDM). The explicit formulations of vehicle dynamics in these CF models make the analyses traceable. However, it is important to be aware that the conclusions rely on the strong assumptions inherent in these models.
@@
A closer look at the JAD-involved simulation reveals how a wave is suppressed by a JAD vehicle: as shown in Figure 5, compared to the queue discharge rate at A, JAD results in a higher flow at B, even though the speed at B cannot exceed that at A due to the constraint imposed by the vehicles at A.
@@
Up until now, according to the latest publication [57], the issue that secondary waves are triggered remains unresolved. The newly triggered secondary waves may largely contribute to increased travel time when absorbing dense stop-and-go waves [54]. Drawing on insights from VSL strategies, we recommend incorporating the empirically observed outflow reduction caused by stop-and-go waves and the potential benefits of mitigating this phenomenon into the modeling process, as it may serve as a key remedy for improving traffic efficiency. As shown in [13], the queue discharge rate reduction is positively and linearly related to the average speed of vehicles within a moving queue, highlighting the potential opportunity of suppressing a wave, especially when the speed within the queue is low.
@@
It is worth noting that we have observed a growing trend in which a platoon of vehicles has been employed and controlled to implement JAD for suppressing stop-and-go waves [58, 59, 60, 56]. However, this approach seems to deviate from the original motivation behind JAD strategies, which is to suppress stop-and-go waves using a single (or very few) vehicle(s). We should be clearly aware that if we have the capability to stably and precisely control a vehicle platoon, the platoon-control-based JAD strategies should be compared with high CAV penetration-based speed homogenization strategies [29, 30, 31, 32, 33, 34], rather than with the single-vehicle-based JAD strategies. In Section III-G, we will illustrate that single-vehicle-based JAD is actually quite feasible in practice.
Another concern with JAD is whether a single JAD vehicle can ensure stable traffic flow behind it. Simply slowing down a JAD vehicle may result in excessively high densities in its wake, potentially triggering a breakdown. In VSL, stabilization is achieved through a combination of controlled speed (enforced by the speed limits) and sufficiently low density (e.g., region CDEF in Figure 3). Therefore, a model of traffic instability urgently needs to be incorporated.
@@ III-A Fundamental Diagrams
Many existing strategies and analyses are based on the FDs or their Lagrangian variants (e.g., speed-spacing relationship). It can be asserted that the shape of the assumed FDs directly influences both the formulation of a strategy and its effectiveness.
As shown in [68], the shape of an FD is closely related to its spatial relation to a bottleneck (Figure 6). Specifically, an FD observed immediately downstream of a bottleneck typically exhibits a clear capacity drop (the second FD from the left in Figure 6), whereas upstream of the bottleneck, the FD tends to approximate a triangular shape (the first FD from the left in Figure 6). Therefore, when assuming an FD, it is crucial to first determine the location or the type of the stop-and-go waves.
@@
For the SPECIALIST, it targets isolated stop-and-go waves (such as the one with the label “A jam wave” in Figure 1(d)). Since the upstream and downstream traffic of such waves are in high-density free-flow conditions, the wave can be regarded as a (moving) bottleneck, and the associated FD may exhibit a capacity drop (Figure 7(b)). This assumed $\lambda$shape FD effectively resolves the contradiction between the widely observed constant length of a moving queue and the possibility of shrinking after imposing a lower speed limit. More specifically, to be consistent with real-world observations (i.e., a constant length of a moving queue), the inflow and outflow of the queue must be equal (i.e., at states 1 and 6 in Figures 7(a) and (b)). Without capacity drop (i.e., no state 5), traffic will return from state 4 to state 1 (illustrated by the blue line and arrow in Figures 7(a) and (b), respectively), leading to the formation of another constant-length queue at state 4. More importantly, this queue at state 4 will not shrink and may continue to propagation upstream, although the density at state 4 is lower than the wave density at state 2. It is noted that state 5 is highly unstable due to high flow [69]. Therefore, even though a stop-and-go wave is suppressed, as shown in Figure 7(a), it remains uncertain how long the high-flow region (state 5) can be sustained.
@@
In contrast, JAD strategies target more broadly defined stop-and-go waves, rather than focusing solely on isolated waves. When multiple waves occur, as shown in the time-space diagrams in Figure 1, the traffic across the multiple and periodic stop-and-go waves is all in the congestion regime, and the corresponding FD is a triangle with no distinct capacity drop (similar to the first FD from the left in Figure 6). Many JAD strategies do not make this distinction clearly, as the original JAD idea is simply to perform “slow-in” and fast-out” actions. For example, state b in Figure 7(c) is still self-sustained [54]. To achieve a self-shrinking JAD-impacted region with the triangular FD, [51] concluded that an effective JAD requires the inflow to be lower than the capacity in the triangular FD (e.g., inflow at state d in Figure 7(d)). It ensures the shrinking of the wave after JAD, but the condition implies that the wave, where JAD can take effect, can also shrink on its own (although it may take time), regardless of whether JAD control is applied (see the green line in Figure 7(c)). More rigorous efforts are called for to establish a robust theoretical foundation grounded in empirical evidence.
Another two assumptions about FDs are also worth mentioning. The first is the use of three-phase FDs in [48, 50], which leads to different conclusions compared to the ones based on the conventional two-phase FDs [44]. Interested readers may look at Figure 6 in [48], which is quite similar to the empirical result shown in Figure 3(a). But we also remind that the TPTT may be explainable by the triangular FD-based KWT and lane-changing (LC) behavior [68]. The second one incorporates the state that is higher than the triangular-shaped states (i.e., state 5 in Figure 7(b)) into the triangle [60, 47] or modifies the shape of the triangular FD to accommodate the higher capacity [41, 70], thereby achieving the goal of simplifying the modeling.
@@
III-B Secondary Waves
There are three types of secondary waves. Type-I secondary waves are the ones that result directly from the implementation of speed reduction of upstream traffic. In SPECIALIST, it corresponds to region EFD in Figure 3(a) and Figure 7(a). In JAD, it is region CEF in Figure 3 of [51]. Even though controlled by VSL, it is still unclear if the traffic in region EFD would breakdown under certain traffic conditions.
Type-II secondary waves appear upstream of the controlled region (e.g., the upstream of line EF in Figure 3(c)). These stop-and-go waves emerge upstream of the speed-limited vehicles, caused by vehicle accumulation at the boundary between a higher-flow upstream region and a lower-flow downstream region. In the SPECIALIST experiment, these new waves resulted from an overly high setting of the target density in region EFD in Figure 3(a) and Figure 7(a). Since this specific traffic state—a particular combination of speed and density—had not previously been realized in practice, the exact stability threshold was unknown. After adjusting to a lower density value, these instabilities largely disappeared.
Unlike VSL, which controls a time-space region, JAD expects to suppress (or say, influence) a time-space region by directly controlling a time-space trajectory. As a result, many existing studies do not clearly distinguish between type-I and type-II secondary waves described above. The prevailing logic in JAD appears to be: dispatch a JAD vehicle, and if it fails to effectively absorb the secondary wave, dispatch another. Consequently, if the wave behind a JAD vehicle diminishes, it is considered that no secondary wave issue exists; if the wave persists or new waves emerge, it is taken as an indication of a secondary wave problem.
Note that most of the secondary wave-related issues mentioned in the previous sections belong to type-II, because a VSL is activated only when the type-I secondary wave is predicted to be finite. Inspired VSL, there is an urgent need to establish clear and reliable criteria for JAD to pre-determine whether activating a JAD vehicle can effectively absorb an upcoming stop-and-go wave. With such criteria, we can make more informed decisions about whether to deploy a JAD vehicle, rather than relying on ad-hoc decisions.
Type-III secondary waves may occur downstream of the controlled region or trajectory, due to the higher outflow resulting from VSL or JAD interventions. However, since stop-and-go waves typically originate from infrastructural bottlenecks with limited capacity, if the intervention induces a forward wave with a flow close to or exceeding the bottleneck capacity, traffic may break down again downstream. A more thorough investigation of Type-III secondary waves is warranted, particularly with respect to their triggering mechanisms, contributing factors, and broader implications.
III-C Generalizability of Strategies